THE X̃^{2}A_{1} − Ã^{2}B_{2} CONICAL INTERSECTION IN NO_{2}, OR HOW EXPERIMENTS AND THEORY IN-TERPLAY TOGETHER

RÉMY JOST, \underline{PATRICK DUPRÉ}, PATRICE THEULÉ, ANTOINE DELON, Grenoble High Magnetic Field Laboratory, CNRS-MPI, BP 166, 25 Rue des Martyrs 38042 GRENOBLE, Cedex 9, France; MARCEL JACON, GSMA, UFR Sciences, BP 1037, 51687 REIMS Cedex 2, France.

We present a comparison between experimental results (vibronic energies, absorption and LIDFS intensity ratios, rotational constants) and ab-initio calculations (diabatic and adiabatic energies, vibronic matrix elements). The three main inputs required in the analysis of the X̃^{2}A_{1} − Ã^{2}B_{2} conical intersection are: i) the diabatic levels of the X̃^{2}A_{1} state (i.e., the complete set of approximatively 200 low lying levels of the X̃^{2}A_{1} observed by LIDFS), ii) the diabatic levels of the Ã^{2}B_{2} state (they are approximately predicted by the ab-initio calculations), iii) the matrix elements of the vibronic interaction, V_{12}, between the X̃^{2}A_{1} and Ã^{2}B_{2} electronic states. The validity of the simplified form proposed for V_{12}, namely “λQ_{3}”, will be discussed. The comparison allows to assign some observed vibronic levels, which in return can be used to improve some parameters of the initial ab-initio PESurfaces. Up to now, our analysis is limited to the four lowest polyads of the Ã^{2}B_{2} state ranging from 9700 cm^{−1} to 12300 cm^{−1}. At higher energy the interactions are stronger, leading to vibronic chaos above ∼ 17000 cm^{−1}.